# RASP: A Drone-based Reconfigurable Actuation and Sensing Platform for Engaging Physical Environments with Foundation Models

### v2

@@ Abstract.
Foundation models and large language models have shown immense human-like understanding and capabilities for generating text and digital media. However, foundation models that can freely sense, interact, and actuate the physical world like in the digital domain is far from being realized. This is due to a number of challenges including: 1) being constrained to the types of static devices and sensors deployed, 2) events often being localized to one part of a large space, and 3) requiring dense and deployments of devices to achieve full coverage. As a critical step towards enabling foundation models to successfully and freely interact with the physical environment, we propose RASP, a modular and reconfigurable sensing and actuation platform that allows drones to autonomously swap onboard sensors and actuators in only $25$ seconds, allowing a single drone to quickly adapt to a diverse range of tasks. We demonstrate through real smart home deployments that RASP enables FMs and LLMs to complete diverse tasks up to $85\%$ more successfully by allowing them to target specific areas with specific sensors and actuators on-the-fly.
@@ 1. Introduction
Foundation models (FM) and artificial general intelligence (AGI) have the potential to bring about massive technological and social change to the physical world by discovering, performing, and continually adapting to new applications and environments without needing to manually embed or program them. While LLMs and FMs have been shown to have human-like understanding and digital media generation capabilities, their ability to respond to events and actuate systems in the physical world are comparatively less explored. There are a number of works that incorporate language models to control physical robotic systems through voice (Vemprala et al., 2023; Nie et al., 2022b), from both industry and academia, such as the Figure01 AGI robot (Figure, 2024). The sensing and control in such scenarios are often localized to the vicinity of the autonomous agent.
Scaling up, there are few works that explore the use of LLMs to actuate our environments, particularly smart homes (King et al., 2023, 2024; Rivkin et al., 2023), where events and actions may occur anywhere in the space. These works generally focus on adapting LLMs as a human-like interface to actuate common internet-connected smart appliances (e.g., speakers, television, air conditioning, etc.). Much like how FMs enable general human language and sensory understanding and responses, our work explores how FMs could be used to enable a diverse range of general interactions with objects and spaces that likely cannot be actuated digitally through code. These interactions form a large portion of our daily interactions.
Figure 1. RASP autonomous payload reconfiguration to execute user specified task
@@
For example, a person may want his/her home to physically bring a snack or medicine; this would require a robotic system that could physically identify and carry the payload. Another person, coming into an office area for the time may want to know the warmest desk to sit at in the building; this would require the building to utilize a dense deployment of temperature sensors. A third person in a chemical lab may want the building to monitor and notify him/her about the results of a chemical reaction that s/he started before leaving; this situation could likely be accomplished with a camera with special capabilities to detect these reactions. While it is possible to enable all these applications, these examples highlight two main challenges that prevent FMs and LLMs for achieving the same amount of autonomy in our physical environments as we have seen in the digital domain:
@@
1. Each new application requires a new device or sensor, often with built-in special capabilities. Each of the three examples we discussed would require the user to purchase or engineer a system to satisfy that task. Evolving new applications and functionality in this way is also not scalable. Another way to state this challenge is that the generality and types of applications available are limited by the functionality of the devices and sensors present.
2. Events, tasks, and actions are often localized. Certain events and tasks a user may require the FM to analyze the entire space, but perform an action at a target location. For example, placing rat poison in areas around the home that will be most effective. If a specific sensing modality needed to make this judgment (e.g., moisture) is not present at a specific location, then the FM cannot obtain the information it needs to carry out the task. This leads to the third challenge.
@@
Towards enabling general task execution by FMs in physical environments, we propose RASP, an adaptive, modular, and reconfigurable actuation and sensing drone platform that allows FMs and LLMs to adapt the same drone to diverse scenarios by efficiently switching out sensors and/or actuators depending on the situation. Figure 1 shows an overview of RASP. The drone lands, communicates with the base station, and specifies the module to swap in. The base station secures the drone and actuates the conveyor belt and gripper to remove and replace the onboard sensor/actuator with the requested module. The unified module interface (UMI) ensures that the sensor/actuator module is firmly attached, connected, and can be accessed by the drone. Finally, the drone flies off with the new module for its new task. RASP is an important step towards FMs and LLMs that can dynamically interact with occupants and the environment and addresses all of the challenges mentioned previously:
@@
To demonstrate the utility of RASP for LLMs and FMs interfacing with physical environments, we prototype and show how RASP could be integrated into a personal assistant system that leverages static sensors deployed throughout the environment (cameras) in conjunction with foundation models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting. While there are several existing configurable drone platforms (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023), these works focus on the physical design and control of the drone rather than the sensing, analysis, and actuation capabilities. Moreover, these existing works require manual reconfiguration by humans, while RASP autonomously reconfigures the sensing and actuation capabilities on-the-fly. To the best of our knowledge, our work is the first to propose a drone platform with reconfigurable sensing to address challenges in coverage, localized sensing, and dense deployments for enabling general LLM and FM interactions with our physical environments. Our contributions are as follows.
1. We propose and implement RASP, a modular and reconfigurable sensing and actuation platform that enables FMs to adapt the same drone to a wide range of scenarios through physical spaces. RASP equips FMs and LLMs with drones that can dynamically change their onboard sensing and actuation modules in under $25$ seconds.
2. To realize RASP, we propose a three-layer architecture that physically replaces onboard modules, maintains data connection lines, and enables drones to communicate with sensing and actuation modules through a unified interface. Unlike existing reconfigurable drone platforms that focus on drone control and design, and require manual intervention to swap components, RASP focuses on systems and methods to autonomously reconfigure the drone’s onboard sensing and actuation capabilities, without human intervention.
3. We demonstrate how RASP can augment LLMs and FMs and easily enable new applications throughout our environments, beyond the capabilities of common IoT smart devices, by prototyping a personal assistant system that leverages both static sensors (cameras) and the mobility and flexibility of a reconfigurable drone. Our deployments in realistic scenarios demonstrate that RASP can improve the success rate of a diverse array of tasks by up to $85\%$.
@@ 2. Related Works
2. Reconfigurable sensing and drone platforms. There are many reconfigurable and modular sensing platforms in existence. In addition to the platforms provided by open-source do-it-yourself (DIY) electronic vendors, such as Adafruit (Industries, 2022) and Sparkfun (Electronics, 2022), there are also platforms that operate on even lower resource microcontrollers, without an operating system, that are less flexible in the number of interfaces and configurations they support (Yi et al., 2018).  (Zhao et al., 2022, 2023) are reconfigurable sensing platforms based on the Raspberry Pi that have unified and generic hardware interfaces, allowing sensors to use the same set of connectors even if they are interfaced differently in software (e.g., UART, SPI, or I2C). Several platforms leverage the Berkeley TinyOS operating system (Levis et al., 2005), which allow developers to develop extremely long-lasting applications with great flexibility (Culler et al., 2002; O’Flynn et al., 2005; Edmonds et al., 2005; Harte et al., 2007; Mikhaylov et al., 2013; Mikhaylov and Huttunen, 2014).
@@
There are a few reconfigurable drone platforms, but most focus on enabling flexible physical design and control of drones (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023). For example, (da Silva Ferreira et al., 2020) allows operators to reconfigure the number of rotors the drone with locking mechanisms that connect rotor modules. These works typically do not focus on sensing and actuation reconfiguration, unlike RASP. Additionally, they typically require a person to manually reconfigure the drone, while we introduce mechanisms that enables human-free configuration of the sensing and actuation. (Xia et al., 2021) recently proposed a reconfigurable drone platform, but is only a demo abstract and leaves out many details on how it autonomously swaps modules.

### v3

# FlexiFly: Interfacing the Physical World with Foundation Models Empowered by Reconfigurable Drone Systems

@@ Abstract.
Foundation models (FM) have shown immense human-like capabilities for generating digital media. However, foundation models that can freely sense, interact, and actuate the physical domain is far from being realized. This is due to 1) requiring dense deployments of sensors to fully cover and analyze large spaces, while 2) events often being localized to small areas, making it difficult for FMs to pinpoint relevant areas of interest relevant to the current task. We propose FlexiFly, a platform that enables FMs to “zoom in” and analyze relevant areas with higher granularity to better understand the physical environment and carry out tasks. FlexiFly accomplishes by introducing 1) a novel image segmentation technique that aids in identifying relevant locations and 2) a modular and reconfigurable sensing and actuation drone platform that FMs can actuate to “zoom in” with relevant sensors and actuators. We demonstrate through real smart home deployments that FlexiFly enables FMs and LLMs to complete diverse tasks up to $85\%$ more successfully. FlexiFly is critical step towards FMs and LLMs that can naturally interface with the physical world.
Embodied AI, Foundation Model Agent, Reconfigurable Drone Platform
[…]
@@ 1. Introduction
While there are a number of works that incorporate large language models (LLM) into robotic and egocentric systems (Vemprala et al., 2023; Nie et al., 2022b), such as the Figure01 AGI robot (Figure, 2024) and digital voice assistants, there are few works that explore the use of LLMs and foundation models (FM) to actuate our physical environments. Unlike egocentric systems, where sensing, processing, and control are often localized to the vicinity of the autonomous agent, executing tasks and monitoring spaces often involves sifting through heterogeneous streams of sensing data spanning large areas of the space to detect and process events localized to a tiny fraction. While FMs and LLMs have shown strong performance on summarization tasks, they have difficulty completing tasks that involve processing localized areas of interest without mechanisms to help them “zoom in”. Much like how FMs enable general human language and sensory understanding and responses, our work explores how FMs could be used to enable a diverse range of general interactions with objects and spaces that likely cannot be actuated digitally through code. These interactions form a large portion of our daily interactions.
Figure 1. FlexiFly enables FMs to “zoom in” to areas of interest with reconfigurable drones to better interface with physical environments.
Figure 2. System architecture of intelligent assistant with FlexiFly.
@@
For example, a person may want his/her home to physically bring a snack or medicine; this would require a robotic system that could physically identify and carry the payload. Another person, coming into an office area for the time may want to know the warmest desk to sit at in the building; this would require the building to utilize a dense deployment of temperature sensors. A third person in a chemical lab may want the building to monitor and notify him/her about the results of a chemical reaction that s/he started before leaving; this situation could likely be accomplished with a camera with special capabilities to detect these reactions. While it is possible to enable all these applications, these examples highlight three main challenges that prevent FMs and LLMs for achieving the same amount of autonomy in our physical environments as we have seen in the digital domain:
@@
1. Each new application requires a new device or sensor, often with built-in special capabilities. Each of the three examples we discussed would require the user to purchase or engineer a system to satisfy that task. Evolving new applications and functionality in this way is also not scalable.
2. Events, tasks, and actions are often localized. Events and tasks are often localized to small areas. As we will show in Section 3, an FM that is analyzing data from multiple sensors covering a large space often misses events occurring in small sections of a large scene. While LLMs and FMs have shown great performance in analyzing general trends, we will show that they have difficulty detecting and responding to localized events in larger sensor deployments (Section 3).
@@
We propose FlexiFly, a platform that addresses the challenges in enabling FMs to interface with the physical world by enabling FMs, particularly visual-language models (VLMs), to “zoom in” to localized areas of interest to obtain a higher resolution, fine-grained, and better understanding of the physical environment to carry out the task at hand. FlexiFly accomplishes this goal with two critical design choices, as shown in Figure 1.
First, we propose a novel image segmentation method called Aspect Ratio Constrained K-Means (ARCK-Means) segmentation to visually “zoom in” and identify potential locations pertinent to the task. We show how existing state-of-art segmentation methods (e.g., Segment Anything Model (SAM) (Kirillov et al., 2023)) often result in split objects that reduce object identification and localization accuracy. ARCK-Means segmentation ensures that full objects are processed and improves the analysis of localized areas of interest in FMs.
Second, we propose an adaptive, modular, and reconfigurable sensing and actuation drone platform that FMs can actuate to “zoom in” physically. Once the FM identifies potential locations of interest, it can select the relevant sensor or actuator to equip, before actuating the drone to analyze locations of interest up close. For example, to answer the question: “where is the warmest place to sit?”, the FM would identify potential locations that may indicate warmth (e.g., sunlight), before sending a drone equipped with a temperature sensor to confirm. While there are several existing configurable drone platforms (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023), these works focus on the physical design and control of the drone rather than the sensing, analysis, and actuation capabilities. Moreover, these existing works require manual reconfiguration, while FlexiFly autonomously reconfigures sensing and actuation capabilities on-the-fly. Our motivation for exploring the use of drones to empower foundation models interfacing with the physical world stems from our vision that humans and robotic systems will coexist in the future. Moreover, our design choices address the challenges mentioned previously:
@@
To demonstrate the utility of FlexiFly for LLMs and FMs interfacing with physical environments, we prototype and show how FlexiFly could be integrated into a personal assistant system that leverages static sensors deployed throughout the environment (cameras) in conjunction with foundation models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting. To the best of our knowledge, our work is the first to propose a drone platform with reconfigurable sensing to address challenges in coverage, localized sensing, and dense deployments for enabling general LLM and FM interactions with our physical environments. Our contributions are as follows.
1. We propose FlexiFly, a platform that enables FMs, particularly VLMs and LLMs, to better understand and interface the physical world by identifying localized areas of interest from large scenes, equipping relevant sensors/actuators, and actuating the drone to “zoom in” up close.
2. To realize FlexiFly, 1) we propose a novel image segmentation method called ARCK-Means that aids FMs in identifying localized areas of interest. We demonstrate that ARCK-Means improves the understanding and detection of objects in large and cluttered scenes over existing segmentation techniques, namely SAM, by reducing the amount of disjoint and split objects produced by segmentation masks. 2) We introduce a drone-based modular and reconfigurable sensing and actuation platform that enables FMs to adapt to a wide range of scenarios. FlexiFly actuates the drone, equipped with task-relevant sensors and actuators, to identify areas of interest to analyze up close.
3. We demonstrate how FlexiFly can augment LLMs and FMs and easily enable new applications throughout our environments, beyond the capabilities of common IoT smart devices, by prototyping a personal assistant system that leverages both static sensors (cameras) and the mobility and flexibility of a reconfigurable drone. Our deployments and demo (Zhao et al., 2024) in realistic scenarios demonstrate that FlexiFly can improve the success rate of a diverse array of tasks by up to $85\%$.
@@ 2. Related Works
2. Reconfigurable and adaptive sensing platforms. There are many reconfigurable and modular sensing platforms in existence. In addition to the platforms provided by open-source do-it-yourself (DIY) electronic vendors, such as Adafruit (Industries, 2022) and Sparkfun (Electronics, 2022), there are also platforms that operate on even lower resource microcontrollers, without an operating system, that are less flexible in the number of interfaces and configurations they support (Yi et al., 2018).  (Zhao et al., 2022, 2023) are reconfigurable sensing platforms based on the Raspberry Pi that have unified and generic hardware interfaces, allowing sensors to use the same set of connectors even if they are interfaced differently in software (e.g., UART, SPI, or I2C). Several platforms leverage the Berkeley TinyOS operating system (Levis et al., 2005), which allow developers to develop extremely long-lasting applications with great flexibility (Culler et al., 2002; O’Flynn et al., 2005; Edmonds et al., 2005; Harte et al., 2007; Mikhaylov et al., 2013; Mikhaylov and Huttunen, 2014).
@@
There are a few reconfigurable drone platforms, but most focus on enabling flexible physical design and control of drones (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023). For example, (da Silva Ferreira et al., 2020) allows operators to reconfigure the number of rotors the drone with locking mechanisms that connect rotor modules. These works typically do not focus on sensing and actuation reconfiguration, unlike FlexiFly. Additionally, they typically require a person to manually reconfigure the drone, while we introduce mechanisms that enables human-free configuration of the sensing and actuation. (Xia et al., 2021) recently proposed a reconfigurable drone platform, but is only a demo abstract and leaves out many details on how it autonomously swaps modules.
Prior works have also explored adaptive sampling strategies to optimize sensor deployment and data collection, both in traditional wireless sensor networks (Lee et al., 2011) and mobile sensing platforms like underwater vehicles (Chen et al., 2012). Unlike these approaches that focus on sampling optimization, FlexiFly enables on-demand sensor reconfiguration guided by foundation models.
3. Embodied systems. Traditional robotic systems often rely on hand-crafted algorithms, making adaptation to new or unforeseen situations challenging. Additionally, they struggle to generalize learned behaviors across different tasks, limiting their effectiveness in real-world applications (Hu et al., 2023). By leveraging LLMs and FMs in robotics, including drones, the generalization problem in planning (Huang et al., 2022; Lin et al., 2023) and control (Chen et al., 2024; Yang et al., 2025; Singh et al., 2022) in different tasks is partially resolved. However, these works still face severe limitations for solving diverse and general tasks due to insufficient real-world interactions. This is primarily because traditional robotic systems rely on limited and fixed sensing capabilities, often restricted to cameras; once a robotic system is built, their hardware capabilities are hardly ever upgraded. Many tasks require multiple sensing modalities or may need the use of different types of sensors to effectively understand, interpret, and respond to events in the physical world. To resolve the aforementioned limitations, we propose FlexiFly, a system that bridges the physical world with foundation models through a modular sensing platform.
3. Preliminaries and Challenges
In this section, we explore the limitations of current foundation models in understanding and responding to events and tasks that require interacting with physical environments. Such tasks a person may ask his/her home or environment could include “where is the warmest place to sit”, “where did I leave my phone”, ”bring me a snack”, and much more.
The growth of the Internet of Things (IoT) is exponential, and the number of internet connected sensors and devices is expected to double by 2030 (Sinha, 2024). It is expected that all future buildings and homes will be equipped with a wide array of different sensors and devices that will improve our efficiency, automation, personalized insights, and overall quality of life. We envision that foundation models will be capable of understanding data collected throughout our physical environments and act as a natural interface between humans and their built environments.
3.1. Deployment
To test the current limitations of FMs for realizing this vision, we created the deployment shown in Figure 8e, using a network of 4 cameras. Since FMs for language and vision are most mature, we focus on these modalities; in the future, we believe that foundation models capable of interpreting streams of heterogeneous sensing modalities from sensors strewn throughout the environment will be possible.
We deploy a camera network on the ceiling of our deployment area, facing directly downwards. We generate floor maps and stitch together frames from individual camera views, according to (Liu et al., 2022), which we input into the Large Language-and-Vision Assistant (LLaVA) (Liu et al., 2023a), a state-of-art visual-language model (VLM), along with user specified tasks and commands.
3.2. Types of Tasks
We identify and explore four classes of tasks in this work that are useful for creating smarter physical spaces, but are not commonly packaged into existing IoT smart devices:
T1: Object/Location Identification (ID). This set of tasks requires the system to observe the environment and identify an object or location based on the user’s command (e.g., “where is my phone?”). In our preliminary analysis, we look at one task in this category: “where is my phone?” and place a phone at different locations in view of the cameras.
T2: Object/Location State (State). This set of tasks involves learning about the state or condition of a specific object or location. For example, “is my food burning” would require the FM to identify food that is being cooked and if the food is producing too much smoke. In our preliminary analysis, we look at one task in this category: “where is the warmest place to sit?” and artificially turn up the heat in certain areas of the room by placing concealed space heaters nearby.
T3: Surveillance (Surv). These tasks differ from the previous two, which are executed only once after the command is received. In surveillance tasks, the system needs to continuously analyze the environment until a potential event occurs (e.g., “Let me know if any chemical in my experiment spills on the table”). In our preliminary analysis, we look at one task in this category: “let me know if any chemical in my experiment spill on the table?” and artificially simulate this event by knocking over a glass of colored water.
T4: Actuation (Act). Unlike previous categories, which require sensing, actuation tasks require direct physical interactions with the environment (e.g., “bring me my medicine”).
3.3. Preliminary Analysis and Limitations
Figure 3. Preliminary study: (a) task completion rate of standard FMs leveraging VLM (camera) and dense sensor networks, compared to FlexiFly; (b) Example showing that the VLM could not detect the phone in plain sight unless “zoomed in”.
Using this basic camera network + FM setup (labeled VLM Baseline), we attempt to satisfy one command from each of the classes of actions we identified. We also compared against an FM analyzing data from a deployed sensor network (labeled Sensor Baseline); here, we deployed 9 sensors (temperature for the state task and alcohol sensors for the surveillance task) in a grid pattern and interpolated between sensors to obtain a map spanning the entire space. We ran 70 trials for each scenario and show the success rate of identifying or completing each task in Figure 3a. We see that the success rate is low across the board for the VLM and sensor baseline, due to three limitations.
Limitation 1: Physical size of environments and volume of sensor information and is large, while events and objects of interest are localized to small areas. A FM overseeing physical environments need to process many streams of data covering a large area, such as the space of our deployment. However, users are generally interested in only a small portion of the environment. We noticed that for the ID tasks (looking for phone) and surveillance (detecting spills), our FM could often not detect these objects and events due to the event occurring in a tiny portion of the environment and the limited resolution that can be captured. As shown in Figure 3b, LLaVA could not identify many objects unless we manually zoom in and reduce the camera’s area of coverage.
Limitation 2: Coverage and resolution of sensing modalities required is limited and not practically scalable. The VLM could not accurately identify the warmest place to sit (T2: state task) because it leverages a sensing modality where standard vision may not perform well (temperature). Even when we included temperature sensors (sensor baseline), the success rate is still low. When we simulated the warmest place to sit close to the sensor, the success rate was high, but the sensor network cannot accurately capture the dynamics of the space at areas away from the sensors, which is where success rate decreased. This highlights that the density of sensors the performance and understanding an FM has about the environment. Additionally, while detecting the warmest place to sit may require a temperature sensor, different applications require different sensor deployments (e.g., detecting falls could use audio or vibration). For instance if audio is used to detect falls and the bedroom does not have a smart speaker or microphone, then it would be difficult to enable this service in the bedroom.
Limitation 3: Actuation is restricted. An actuation task such as bringing the user a snack or medicine cannot be accomplished with only an FM along and static devices throughout the environments.
3.4. Design Philosophy
To address the limitations discussed previously, our design philosophy involves creating mechanisms that enable FMs to “zoom in” and sense targeted areas of interest with higher resolution. We tackle these limitations on two fronts. First, we propose a novel image segmentation technique called Aspect Ratio Contrained K-Means (ARCK-Means) to aid in identifying potential locations of interest to “zoom in”. Second, we propose FlexiFly, a drone platform, with modular and reconfigurable sensing and actuation, that allows FMs to pick and choose sensors and actuators depending on the task at hand. The FM first leverages ARCK-Means to identify areas of interest based on the current state of the environment and the input command (Section 4), before actuating the drone with the corresponding sensors to “zoom in”(Section 5).
For example, an FM looking to answer “where is the warmest place to sit” previously could only rely on analyzing and sifting through a large amount of sensor data from static sensors throughout the environment. With FlexiFly, an FM identifies local areas of interest with ARCK-Means, before actuating a temperature sensor equipped drone to areas of interest to measure and compare temperature readings.
ARCK-Means alleviates the first limitation by aiding the FM in identifying local areas of interest in large cluttered scenes. Our reconfigurable sensing and actuation drone platform alleviates limitation 2 and 3. Instead of requiring a dense deployment of static sensors (limitation 2), FMs can reconfigure and actuate the drone to the precise location it needs to sense, which reduces deployment overhead. Moreover, a mobile drone platform enables actuation of items in the physical environment that are more restrictive for static sensors and devices (limitation 3).
4. Identifying Local Areas of Interest via Segmentation.
Figure 4. Segmentation and clustering to break down scenes into smaller more manageable pieces for LLaVA and DINO. (a) Object masks after applying Segment Anything Model (SAM); (b) Extracted frames after clustering object masks based on K-Means; (c) Hierarchical clustering, and (d) ARCK-Means. For ARCK-Means, we constrain the aspect ratio of extracted frames to be between 0.67 and 1.5.
As mentioned in Section 3, when a user command arrives, we input the command along with frames stitched together from the camera network in LLaVA. We found that LLaVA was not able to reliably detect objects of interest in a scene when we passed in an image or a large space (e.g., a single chair in a large room); the larger the scene, the less detailed and more high-level the responses became.
One way to alleviate this is to segment and analyze smaller portions of the full image. Instead of directly passing the entire scene into LLaVA, we use the state-of-art image segmentation model, Segment Anything Model (SAM), to extract key objects and smaller areas of the scene (Kirillov et al., 2023). Figure 4a shows an example of the object masks that are output by SAM. These object masks are then input into LLaVA, along with the user command, and outputs (in text) the names of potential objects in the scene that might be of interest. These text outputs from LLaVA and the segmented images are input into Grounding DINO (Liu et al., 2023b), a state-of-art language model for zero-shot object detection and localization, which then outputs localized bounding boxes of potential objects of interest. In implementing this segmentation pipeline for identifying locations of interest, we noticed a number of limitations that impacted performance, as discussed next.
4.1. Limitations of Current Segmentation Methods
1. SAM often generates segmentation masks that split objects between two different masks. As such, LLaVA and Grounding DINO often identify the same object multiple times at slightly different locations when directly using the masks from SAM, adding to the processing time.
2. The segmented masks are not rectangular, which reduces object detection performance. We observed that directly inputting the segmented masks into LLaVA and Grounding DINO saw a large performance degradation. This performance degradation persisted even if we zero pad the segmented images into a rectangular shape. We suspect this is because LLaVA and Grounding DINO are typically trained with rectangular images, with standard aspect ratios (e.g., 640 x 480), in natural settings; whereas, the inputs we were trying are non-rectangular with most of the background removed (zero padded).
4.2. Aspect Ratio Constrained K-Means (ARCK-Means) Segmentation
To address the limitations of current segmentation methods in aiding FMs identify potential locations of interest, we propose Aspect Ratio Constrained K-Means (ARCK-Means) Segmentation. ARCK-Means operates on the segmented objects generated by SAM.
To alleviate the challenge of splitting objects between different masks (limitation 1), ARCK-Means first clusters the centroids of the masks generated into $k$ clusters. We leverage hierarchical clustering (Murtagh and Contreras, 2012), but also benchmark against K-means (Ahmed et al., 2020).
To ensure that whole rectangular images are analyzed by LLaVA and Grounding DINO (limitation 2), we find the minimum-area rectangle that encompasses all masks of each cluster. Moreover, we constrain the aspect ratio of clustered masks between to ensure aspect ratios of typical images used during training, resulting in the full ARCK-Means method. To accomplish this, we check if adding a new segment into the existing cluster causes the aspect ratio to fall below or above our constraints; if these constraints are broken, then we do not make the assignment. The resulting image segments are then analyzed by LLaVA and DINO.
4.3. Analysis and Benchmarking
First, we tested two prompts with the LLaVA model and vary the size of the object with respect to the frame of one camera (Figure 5a). The first prompt is more general, asking LLaVA to describe the objects present. The second prompt is more specific, asking if the specific object of interest is present in the scene. We see that the more specific a prompt is, the higher the recall, which is the reason why we used more specific prompting. Second, we see that as the size of the object gets smaller with respect to the image frame, the recall gets smaller, since the signal-to-noise ratio of the object gets smaller.
In Figure 5b, we compare the recall of each clustering method after clustering each scene into 5 segments, and see that ARCK-Means has the highest recall due to improvements in the masks it clusters (addressing limitation 1) and the shape of the resulting segments (addressing limitation 2); as such, we adopt ARCK-Means into the final system. Figure 5c shows the run time and recall of ARCK-Means as a function of the number of clusters or segments we split the scene into. We see that the recall levels out at around $90\%$ after $5$ clusters, while taking around $100$ms to run the full pipeline. Adding in more clusters does not yield significant improvements in detection, but significantly increases run time; as such, we segment the scene into five segments in the final system. To generate these plots, we used 47 images of indoor home, office, and lab environments. We took 35 images from the ADE20K scene parsing dataset (Zhou et al., 2017) as well as 12 images from our own deployment. We implement and run the full visual-language model pipeline (SAM, clustering, LLaVA, and DINO) on an Nvidia RTX 3090 GPU server.